# Bayesian Model Averaging in Causal Instrumental Variable Models

### v4

@@ Abstract
Instrumental variables are a popular tool to infer causal effects under unobserved confounding, but choosing suitable instruments is challenging in practice. We propose gIVBMA, a Bayesian model averaging procedure that addresses this challenge by averaging across different sets of instrumental variables and covariates in a structural equation model. This allows for data-driven selection of valid instruments and provides additional robustness against invalid instruments. Our approach extends previous work through a scale-invariant prior structure and accommodates non-Gaussian outcomes and treatments, offering greater flexibility than existing methods. The computational strategy uses conditional Bayes factors to update models separately for the outcome and treatments. We prove that this model selection procedure is consistent. In simulation experiments, gIVBMA outperforms current state-of-the-art methods. We demonstrate its usefulness in two empirical applications: the effects of malaria and institutions on income per capita and the returns to schooling. A software implementation of gIVBMA is available in Julia.
Keywords: Conditional Bayes factors, Endogeneity, Invalid instruments, Many instruments, Unobserved confounding
@@ 1 Introduction
Instrumental variables offer a way to infer causal effects in the presence of unobserved confounding. To be suitable, an instrumental variable must meet two criteria: It must not be affected by the unobserved confounder and must not affect the outcome directly (valid), and it must be associated with the regressor of interest (relevant). In practice, finding variables that fulfill these assumptions is challenging. Furthermore, when multiple instrumental variables are available, the results can be highly sensitive to the specific instruments chosen. In many applications, there is a large degree of uncertainty over which instruments to use in addition to the usual uncertainty over covariate inclusion in a regression model. To address this uncertainty, we propose a Bayesian model averaging (BMA) procedure that averages across different sets of instrumental variables and covariates in a multivariate sampling model with correlated residuals. In our context, endogenous variables are treatments that are affected by unobserved confounding.
@@
An important contribution of this paper is a data-driven approach for selecting valid instruments. The user specifies a matrix of potential instruments and covariates, and the BMA framework identifies relevant and valid instruments as those variables included in the treatment model but excluded from the outcome model. This approach provides robustness against the use of invalid or irrelevant instruments. We establish a consistency result showing that the posterior asymptotically concentrates on the correct model.
We build on earlier work by Karl and Lenkoski, (2012). In particular, our computational strategy is closely inspired by theirs: we use a Gibbs structure to update the outcome, treatment, and covariance parameters separately and use conditional Bayes factors in the model updates. However, we make a number of contributions that go beyond Karl and Lenkoski, (2012):
  1. 1.
    A careful analysis of prior structures is provided, and we use suitably adapted $g$-priors on the regression coefficients, which make the analysis scale-invariant. Our approach uses independent priors on the outcome and treatment model space without enforcing identification through zero prior probability on non-identified models.
  2. 2.
    We adapt the latent Gaussian framework with univariate link functions (ULLGM) proposed in Steel and Zens, (2025) to allow for non-Gaussian outcomes and endogenous variables. This greatly extends the applicability of our method while not substantially adding to the computational cost.
  3. 3.
    We prove that our model selection procedure is consistent in the sense that the conditional Bayes factors used in the model updates tend to infinity in favor of the true model as the sample size grows.
  4. 4.
    The method proposed in this paper is implemented in the gIVBMA.jl package written in the Julia language (Bezanson et al.,, 2017). While Karl and Lenkoski, (2012) previously released an R package for their method, it has since been withdrawn from CRAN.
A subsequent paper by Lee and Lenkoski, (2022) addresses multiple endogenous and non-Gaussian variables. However, they only focus on the Poisson case, which they analyse using a Laplace approximation. Through the ULLGM framework, we provide a flexible and exact way of incorporating a wide range of non-Gaussian variables, which can easily be adapted to the data. We refer to the method of Karl and Lenkoski, (2012) and Lee and Lenkoski, (2022) as IVBMA and call our proposed procedure generalised IVBMA (gIVBMA).
Through extensive experiments on real and simulated data, we evaluate gIVBMA against IVBMA, BMA, and a number of classical methods tailored for this problem. The gIVBMA methodology works well, especially in settings with potentially invalid instruments. Our proposed procedure also performs well when many individually weak instruments are available. Another advantage of gIVBMA (and IVBMA) over most classical methods is that we obtain the posterior distribution of the structural covariance matrix which informs us on the degree of endogeneity. This also provides these methods with the ability to borrow strength from the model for the endogenous variables.
After a brief review of related literature, Section 2 introduces the sampling model and discusses model uncertainty and identification. Sections 3 and 4 respectively describe the prior specification and posterior inference, while Section 5 provides the model consistency results. Simulated and real data are analysed in Sections 6 and 7. Finally, Section 8 concludes.
1.1 Related literature
@@
Rossi et al., (2005) propose a simple Bayesian approach to instrumental variables (IV) based on conjugate normal and inverse Wishart priors and a Gibbs sampler iterating between the outcome, treatment, and covariance parameters. Koop et al., (2012) were among the first to propose BMA in IV models, offering a general approach for simultaneous equations. However, their complex reversible jump MCMC scheme (which can mix quite slowly in practice) makes their method less accessible for practitioners. Karl and Lenkoski, (2012) build on the Bayesian IV framework of Rossi et al., (2005), using conditional Bayes factors to iteratively update the outcome and treatment models within a Gibbs sampler. Lee and Lenkoski, (2022) further extend this approach to approximately accommodate non-Gaussian (Poisson) distributions. Lenkoski et al., (2014) introduce a hybrid Bayesian version of the classical Two-Stage Least Squares (TSLS) estimator.
@@
Our proposed method also outperforms both IVBMA and state-of-the-art classical estimators in cases with many individually weak instruments. Traditional classical estimators tend to exhibit substantial bias and can even become inconsistent in such settings. To address these issues, common approaches include first-stage regularisation [^1] Inspired by the TSLS procedure, “first-stage” refers to the estimation of the treatment model, while “second-stage” relates to inference in the outcome model. (e.g. Okui,, 2011; Belloni et al.,, 2012; Carrasco,, 2012), using jackknife-fitted values in the first stage (e.g. Angrist et al.,, 1999; Hansen and Kozbur,, 2014), selecting instruments based on minimizing a mean-square error criterion (e.g. Donald and Newey,, 2001), or model-averaging to obtain model-averaged first-stage predictions of the endogenous variable (e.g. Kuersteiner and Okui,, 2010).
1.2 Notation
Let $p(\bm{x})$ denote the probability density function (pdf) or probability mass function (pmf) of the random variable $\bm{x}$. We use $\bm{x}\sim N(\bm{\mu},\bm{\Sigma})$ to indicate that the random vector $\bm{x}\in\Re^{k}$ follows a Gaussian distribution with mean vector $\bm{\mu}\in\Re^{k}$ and positive definite symmetric (PDS) $k\times k$ covariance matrix $\bm{\Sigma}$. We say an $n\times k$ matrix $\bm{R}$ follows a matrix normal distribution, $\bm{R}\sim MN(\bm{M},\bm{S},\bm{T})$ if and only if $vec(\bm{R})\sim N(vec(\bm{M}),\bm{T}\otimes\bm{S}),$ where $\bm{M}\in\Re^{n\times k}$ and $\bm{S}$ and $\bm{T}$ are PDS matrices, respectively of dimension $n\times n$ and $k\times k$.
@@ 2 The Sampling Model
We consider a structural model with a single outcome and $l\geq 1$ endogenous variables:
@@
We allow all instruments to enter the outcome model, which is essential for our model selection procedure to identify valid instruments in a data-driven manner. For valid instruments, $\bm{\beta}$ will have most of its posterior mass on (or very close to) zero, but we do not enforce this a priori. This point will become clearer in the next section, where we introduce model uncertainty.
@@
We assume the residuals are jointly normal $\bm{[\epsilon:H]}\sim MN(\bm{0},\bm{I_{n}},\bm{\Sigma})$, where the structural covariance matrix $\bm{\Sigma}$ can be partitioned into
@@
where $\sigma_{y|x}=\sigma_{yy}-\bm{\Sigma_{yx}\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal}$ and $\bm{H}=\bm{X}-(\bm{\iota}\bm{\Gamma}+\bm{Z}\bm{\Delta})$. The marginal distribution of the treatment matrix, $\bm{X}\sim MN(\bm{\iota\Gamma+Z\Delta},\bm{I_{n}},\bm{\Sigma_{xx}}),$ then completes the joint distribution of $\bm{y}$ and $\bm{X}$.
@@ 2.3 Non-Gaussian models
$F^{(y)}$ and $F^{(x_{j})}$ are the distributions of $\bm{y}$ and $\bm{x_{j}}$, and $h_{y}$ and $h_{x_{j}}$ are invertible univariate link functions that map the latent Gaussian to the appropriate parameter of $F^{(y)}$ and $F^{(x_{j})},j=1\dots,l$. If required, $r_{y}$ and $r_{x_{j}}$ group any additional parameters of $F^{(y)}$ and $F^{(x_{j})}$, which are assumed to be the same for all observations. We also assume the endogenous variables are independent conditional on their latent Gaussian representation, that is, all the dependence is in the Gaussian part. Many distributions can be expressed as members of the ULLGM family. Two examples that we use in our simulations and applications are (for more examples see Steel and Zens,, 2025):
@@ 3.1 Regression coefficients
where $\bm{V}_{M}=[\bm{\iota:Z}_{M}]$ and the other elements of $\bm{\Lambda}$ are zero. We will assume that for all models that we consider, the matrices $\bm{U}_{L}$ and $\bm{V}_{M}$ are of full column rank. This will ensure that the inverses in (2) and (3) exist for all models. Since the $g$-priors include the intercept, we recommend centering all the Gaussian components of $\bm{y}:\bm{X}]$ to ensure that the zero mean prior on the intercept is reasonable.
These priors are conditionally conjugate for our sampling model, so we obtain closed-form expressions for conditional posteriors and marginal likelihoods (see Subsection 4.2). In addition, we only have to elicit two scalar hyperparameters, $g_{L}$ and $g_{M}$. The choice of $g_{L},g_{M}>0$ controls the prior variance and the complexity penalty of the Bayes factors.
We can either fix $g_{L}$ and $g_{M}$ or put hyperpriors on them. We will mostly focus on two possible choices: an adapted version of the benchmark or BRIC prior (Fernández et al.,, 2001) and the hyper-$g/n$ prior (Liang et al.,, 2008). In the adapted benchmark prior, we fix $g_{L}=\max\left\{n,(k+l+1)^{2}\right\}$ and $g_{M}=\max\left\{n,(k+p+1)^{2}\right\}$. In the context of the standard normal linear regression model, the benchmark prior results in model selection consistency (Fernández et al.,, 2001), which will be examined in our setting in Subsection 5.
@@
3.2 Covariance Matrix
As in Karl and Lenkoski, (2012), we put an inverse Wishart prior on the structural covariance matrix, i.e. $\bm{\Sigma}\sim W^{-1}(\nu,\bm{I_{l+1}})$. Centering the prior over the (diagonal) identity matrix reflects the fact that we want inference on the degree of endogeneity to be primarily driven by the data. The degrees of freedom parameter $\nu$ controls the amount of information in the prior and must satisfy $\nu>l$ so that the prior is proper. The choice of $\nu$ can be quite influential as i.a. it controls how tight the prior on all off-diagonal elements of $\bm{\Sigma}$ is around zero. It can substantially impact the inference on the marginal variances and the degree of endogeneity. To avoid being too dogmatic, we assume a hyperprior on $\nu$. An Exponential prior that is shifted by at least $l$, such that $\nu>l$, works well in our experience. For more numerical stability, it may be desirable to shift it slightly further. We use an Exponential with scale $1$ shifted by $l+1$ as our default choice throughout the remainder of the paper. This prior is quite uninformative on $\nu$ itself while implying a prior on the covariance ratio $\bm{\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal}$ that is (almost) identical to fixing $\nu=3$ (as done by Karl and Lenkoski,, 2012). Figure S.1 in the Supplementary Material illustrates the implied priors on the covariance ratio and the outcome variance for shifted Exponential priors with different scales.
3.3 Model Space

### v5

# arXiv:2504.13520v5[stat.ME] 27 Feb 2026

@@ Abstract
Instrumental variables are a popular tool to infer causal effects under unobserved confounding, but choosing suitable instruments is challenging in practice. We propose gIVBMA, a Bayesian model averaging procedure that addresses this challenge by averaging across different sets of instrumental variables and covariates in a structural equation model. This allows for data-driven selection of valid and relevant instruments and provides additional robustness against invalid instruments. Our approach extends previous work through a scale-invariant prior structure and accommodates non-Gaussian outcomes and treatments, offering greater flexibility than existing methods. The computational strategy uses conditional Bayes factors to update models separately for the outcome and treatments. We prove that this model selection procedure is consistent. In simulation experiments, gIVBMA outperforms current state-of-the-art methods. We demonstrate its usefulness in two empirical applications: the effects of malaria and institutions on income per capita and the returns to schooling. A software implementation of gIVBMA is available in Julia.
Keywords: Conditional Bayes factors, Endogeneity, Invalid instruments, Identification, Unobserved confounding
@@ 1 Introduction
Instrumental variables (IV) offer a way to infer causal effects in the presence of unobserved confounding. To be suitable, an instrumental variable must meet two criteria: It must not be affected by the unobserved confounder and must not affect the outcome directly (valid), and it must be associated with the regressor of interest (relevant). In practice, finding variables that fulfill these assumptions is challenging. Furthermore, when multiple instrumental variables are available, the results can be highly sensitive to the specific instruments chosen. In many applications, there is a large degree of uncertainty over which instruments to use in addition to the usual uncertainty over covariate inclusion in a regression model. To address this uncertainty, we propose a Bayesian model averaging (BMA) procedure that averages across different sets of instrumental variables and covariates in a multivariate sampling model with correlated residuals. In our context, endogenous variables are treatments that are affected by unobserved confounding.
@@
An important contribution of this paper is a data-driven approach for selecting valid and relevant instruments. The user specifies a matrix of potential instruments and covariates, and the BMA framework identifies valid and relevant instruments as those variables included in the treatment model but excluded from the outcome model. This approach provides robustness against the use of invalid or irrelevant instruments. We establish a consistency result showing that the posterior asymptotically concentrates on the correct model.
@@
We build on a large existing literature. Key ideas and methodology for Bayesian variable selection in multivariate regression models can be found in Brown et al., (1998), while Rossi et al., (2005) propose a simple Bayesian approach to instrumental variables based on conjugate normal and inverse Wishart priors and a Gibbs sampler iterating between the outcome, treatment, and covariance parameters. Koop et al., (2012) were among the first to propose BMA in IV models, offering a general approach for simultaneous equations. However, their complex reversible jump MCMC scheme (which can mix quite slowly in practice) makes their method less accessible for practitioners. Karl and Lenkoski, (2012) build on the Bayesian IV framework of Rossi et al., (2005), using conditional Bayes factors to iteratively update the outcome and treatment models within a Gibbs sampler. Subsequent work (Lenkoski and Aanes,, 2020; Kourtellos et al.,, 2020; Lee and Lenkoski,, 2022) extends their approach to allow for non-Gaussian and multiple endogenous variables. Lenkoski et al., (2014) introduce a hybrid Bayesian version of the classical Two-Stage Least Squares (TSLS) estimator.
Our proposed methodology follows Karl and Lenkoski, (2012), and we take close inspiration from their computational strategy: we use a Gibbs sampler to update the outcome, treatment, and covariance parameters separately and use conditional Bayes factors in the model updates. However, we also provide a careful analysis of prior structures, and we use suitably adapted $g$-priors on the regression coefficients, which make the analysis scale-invariant, and a flexible Cholesky-based prior on the structural covariance (Lopes and Polson,, 2014). Our approach uses independent priors on the outcome and treatment model space without enforcing identification through zero prior probability on non-identified models. We accommodate non-Gaussian outcomes and endogenous variables by adapting the latent Gaussian framework with univariate link functions (ULLGM) proposed in Steel and Zens, (2026). This greatly extends the applicability of our method while not substantially adding to the computational cost. We prove that our model selection procedure is consistent in the sense that the conditional Bayes factors used in the model updates tend to infinity in favor of the true model as the sample size grows. Finally, we provide software in the accompanying gIVBMA.jl package written in the Julia language. While Karl and Lenkoski, (2012) previously released an R package for their method, it has since been withdrawn from CRAN. We refer to the method of Karl and Lenkoski, (2012) as IVBMA and call our proposed procedure generalised IVBMA (gIVBMA).
Related ideas and methods can also be found in the context of Bayesian inference on binary treatment effects under endogenous treatment selection. Methodology for these models with Bayesian variable selection was developed by Jacobi et al., (2016) using a single latent factor, and by Wagner et al., (2023) from a bi-factor model perspective.
@@
Our proposed method also performs well in cases with many individually weak instruments. Traditional classical estimators tend to exhibit substantial bias and can even become inconsistent in such settings. To address these issues, common approaches include first-stage regularisation [^1] Inspired by the TSLS procedure, “first-stage” refers to the estimation of the treatment model, while “second-stage” relates to inference in the outcome model. (e.g. Okui,, 2011; Belloni et al.,, 2012; Carrasco,, 2012), using jackknife-fitted values in the first stage (e.g. Angrist et al.,, 1999; Hansen and Kozbur,, 2014), selecting instruments based on minimizing a mean-square error criterion (e.g. Donald and Newey,, 2001), or model-averaging to obtain model-averaged first-stage predictions of the endogenous variable (e.g. Kuersteiner and Okui,, 2010).
Through extensive experiments on real and simulated data, we evaluate gIVBMA against IVBMA, BMA, and a number of classical methods tailored for this problem. The gIVBMA methodology works well in settings with potentially invalid instruments and many individually weak instruments. Another advantage of gIVBMA (and IVBMA) over most classical methods is that we obtain the posterior distribution of the structural covariance matrix which informs us on the degree of endogeneity. This also provides these methods with the ability to borrow strength from the model for the endogenous variables.
Section 2 introduces the sampling model and discusses model uncertainty and identification. Sections 3 and 4 respectively describe the prior specification and posterior inference, while Section 5 provides the model consistency results. Simulated and real data are analysed in Sections 6 and 7. Finally, Section 8 concludes.
@@ 2 The sampling model
2.1 Model Specification
We consider a structural model for $n$ observations of a single outcome and $l\geq 1$ potentially endogenous variables:
@@ 2.1 Model Specification
We allow all instruments to enter the outcome model, which is essential for our model selection procedure to identify valid and relevant instruments in a data-driven manner. For valid instruments, $\bm{\beta}$ will have most of its posterior mass on (or very close to) zero, but we do not enforce this a priori. This point will become clearer in the next section, where we introduce model uncertainty.
@@
We assume the residuals follow a matrix normal distribution (defined in supplementary Section S.1) $\bm{[\epsilon:H]}\sim MN(\bm{0},\bm{I_{n}},\bm{\Sigma})$, where the structural covariance matrix $\bm{\Sigma}$ can be partitioned into
@@
where $\sigma_{y|x}=\sigma_{yy}-\bm{\Sigma_{yx}\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal}$ and $\bm{H}=\bm{X}-(\bm{\iota}\bm{\Gamma}+\bm{Z}\bm{\Delta})$. The marginal distribution of the treatment matrix, $\bm{X}\sim MN(\bm{\iota\Gamma+Z\Delta},\bm{I_{n}},\bm{\Sigma_{xx}})$ completes the joint distribution of $\bm{y}$ and $\bm{X}$.
@@ 2.4 Non-Gaussian models
$F^{(y)}$ and $F^{(x_{j})}$ are the distributions of $\bm{y}$ and $\bm{x_{j}}$, and $h_{y}$ and $h_{x_{j}}$ are invertible univariate link functions that map the latent Gaussian to the appropriate parameter of $F^{(y)}$ and $F^{(x_{j})},j=1\dots,l$. If required, $r_{y}$ and $r_{x_{j}}$ group any additional parameters of $F^{(y)}$ and $F^{(x_{j})}$, which are assumed to be the same for all observations. We also assume the endogenous variables are independent conditional on their latent Gaussian representation, that is, all the dependence is in the Gaussian part. Many distributions can be expressed as members of the ULLGM family. Two examples that we use in our simulations and applications are (for more examples see Steel and Zens,, 2026):
@@ 3.1 The prior on the regression coefficients
where $\bm{V}_{M}=[\bm{\iota:Z}_{M}]$ and the other elements of $\bm{\Lambda}$ are zero. We will assume that for all models that we consider, the matrices $\bm{U}_{L}$ and $\bm{V}_{M}$ are of full column rank. This will ensure that the inverses in (2) and (3) exist for all models. Since the $g$-priors include the intercept, we recommend centering all the Gaussian components of $\bm{y}:\bm{X}]$ to ensure that the zero mean prior on the intercept is reasonable. As stated in Subsection 2.2, we impose the same zero restrictions on all treatment model equations if $l>1$. This means that the matrix normal prior in (3) can be adopted, but in case we would want to allow potentially different covariates in each treatment model equation, we would require more flexible priors such as e.g. the Recursive Extended Natural Conjugate prior of Richard and Steel, (1988). In case we would want to impose asymmetric shrinkage across treatment equations, we could adopt the Asymmetric Conjugate prior introduced by Chan, (2022) in the context of VARs. Of course, the size of the model space and the complexity of the analysis would increase substantially and, for our purposes in this paper, we are typically not really interested in the coefficients in the treatment model. Hahn et al., (2018) propose a factor shrinkage prior for the instrument coefficients when $l=1$, based on an approximation to the horseshoe prior (see Subsection S.5.2).
The priors in (2) and (3) are conditionally conjugate for our sampling model, so we obtain closed-form expressions for conditional posteriors and marginal likelihoods (see Subsection 4.2). In addition, we only have to elicit two scalar hyperparameters, $g_{L}$ and $g_{M}$. The choice of $g_{L},g_{M}>0$ controls the prior variance and the complexity penalty of the Bayes factors.
We can either fix $g_{L}$ and $g_{M}$ or put hyperpriors on them. We will focus on two possible choices: an adapted version of the benchmark or BRIC prior (Fernández et al.,, 2001) and the hyper-$g/n$ prior (Liang et al.,, 2008). In the adapted benchmark prior, we fix $g_{L}=\max\left\{n,(p+l+1)^{2}\right\}$ and $g_{M}=\max\left\{n,(p+1)^{2}\right\}$. In the context of the standard normal linear regression model, the benchmark prior results in model selection consistency (Fernández et al.,, 2001), which will be examined in our setting in Subsection 5.
@@
3.2 The prior on the covariance matrix
As in Karl and Lenkoski, (2012), we first put an inverse Wishart prior (as defined in supplementary Section S.1) on the structural covariance matrix, $\bm{\Sigma}\sim IW(\nu,\bm{I_{l+1}})$. Centering the prior over the (diagonal) identity matrix reflects the fact that we want inference on the degree of endogeneity to be primarily driven by the data. We recommend standardising the outcome and endogenous variables (or at least selecting appropriate units of measurement) to ensure the prior scale is not in strong conflict with the likelihood. The degrees of freedom parameter $\nu$ controls the amount of information in the prior and must satisfy $\nu>l$ so that the prior is proper. The choice of $\nu$ can be quite influential as i.a. it controls how tight the prior on all off-diagonal elements of $\bm{\Sigma}$ is around zero. It can substantially impact the inference on the marginal variances and the degree of endogeneity. To avoid being too dogmatic, we assume a hyperprior on $\nu$. An Exponential prior that is shifted by at least $l$, such that $\nu>l$, works well in our experience. For more numerical stability, it may be desirable to shift it slightly further. We use an Exponential with scale $1$ shifted by $l+1$ as our default choice throughout the remainder of the paper. This prior is quite uninformative on $\nu$ itself while implying a prior on the covariance ratio $\bm{\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal}$ that is (almost) identical to fixing $\nu=3$ (as done by Karl and Lenkoski,, 2012). Figure S.1 in the supplementary Subsection S.3.1 illustrates the implied priors on the covariance ratio and the outcome variance for shifted Exponential priors with different scales.
As a more flexible alternative, we also consider the Cholesky-based prior proposed in Lopes and Polson, (2014). Write $\bm{\Sigma}$ as
$\bm{\Sigma}=\begin{pmatrix}1&\bm{a_{yx}}\\ 0&\bm{I_{l}}\end{pmatrix}\begin{pmatrix}\sigma_{y\mid x}&0\\ 0&\bm{\Sigma_{xx}}\end{pmatrix}\begin{pmatrix}1&0\\ \bm{a_{yx}}^{\intercal}&\bm{I_{l}}\end{pmatrix}$, (4)
where $\bm{a_{yx}}=\bm{\Sigma_{yx}}\bm{\Sigma_{xx}}^{-1}$ denotes the scaled residual covariance. We specify the prior on $\bm{\Sigma}$ in three parts by assigning priors to the conditional outcome variance $\sigma_{y\mid x}$, the scaled covariance $\bm{a_{yx}}$, and the treatment variance $\bm{\Sigma_{xx}}$. This allows us to separately quantify the prior uncertainty about these individual covariance components, a potential advantage over the inverse Wishart prior used in Karl and Lenkoski, (2012).
A convenient choice is a Gaussian prior on the scaled covariance, $\bm{a_{yx}}^{\intercal}\sim N(0,\bm{\Omega}_{a})$, an inverse Gamma prior (as in Section S.1) on the outcome variance, $\sigma_{y\mid x}\sim IG(c,d)$, and an inverse Wishart prior on the treatment covariance matrix, $\bm{\Sigma_{xx}}\sim IW(\xi,\bm{I_{l}})$. These priors lead to conditional closed-form updates. As default choices for the variance hyperparameters, we use $c=\nu/2,d=1/2,\xi=\nu-1$, where $\nu$ denotes the degrees-of-freedom parameter of the full inverse Wishart prior. Again, it can be useful to put a hyperprior on $\nu$. For the (scaled) residual covariance, we set $\bm{\Omega_{a}}=\omega_{a}\bm{I_{l}}$, where higher $\omega_{a}$ indicates a weaker prior belief in exogeneity. The key difference with the inverse Wishart prior is that we can now freely choose $\omega_{a}$, whereas it is restricted to $\omega_{a}=\sigma_{y\mid x}$ in the prior implied by the full inverse Wishart on $\bm{\Sigma}$.
3.3 The prior on the model space